OpenStudy (anonymous):

Solve. Please show the algebraic inequality you used and show all of your work. Two cell phone companies are advertising rates. Company A charges a rate of $20 per month plus $0.05 per minute. Company B charges a rate of $10 per month plus $0.10 per minute. What is the number of minutes used above which Company A costs more than Company B?

OpenStudy (anonymous):

20 + x(.05)=10 + x(.1) -10 both sides 10 + x(.05) = x(.1) - x(.05) both sides 10 = .05x divide by .05 both sides x= 200 i think its been a while

OpenStudy (anonymous):

you are my hero lol

OpenStudy (anonymous):

if I a m right